OpenStudy (anonymous):

Maximize x^2+y^2 subject to the condition ax^2+2hxy+by^2=1.

OpenStudy (freckles):

\[P=x^2+y^2 \\ ax^2+2hxy+by^2=1 \\a \cdot x^2+2hy \cdot x+by^2-1=0 \\ \ \text{ Solve the second equation for x} \\ x=\frac{-a \pm \sqrt{(2hy)^2-4(a)(by^2-1)}}{2(a)}\] That's really ugly... So you can replace x with that and differentiate P w.r.t. y

OpenStudy (freckles):

well -b would be -2hy

OpenStudy (freckles):

\[x=\frac{-\color{red}{2hy} \pm \sqrt{(2hy)^2-4(a)(by^2-1)}}{2(a)} \]

OpenStudy (paxpolaris):

** correction:\[x=\frac{-\color{red}{2hy} \pm \sqrt{(2hy)^2-4(a)(by^2-1)}}{2(a)}\] \[\implies x={-hy \pm \sqrt{\left( h^2-ab \right) y^2+a}\over a}\]

OpenStudy (freckles):

By the way pax and I are just using the quadratic formula. I know it looks scarier than that.

OpenStudy (anonymous):

There is Lagrange's method to solve , which may be easier I think, but its huge.

OpenStudy (freckles):

\[\text{ Let } P=x^2+y^2 \\ \text{ and } G=ax^2+2hxy+by^2 \\ P_x=\lambda G_x \\ P_y=\lambda G_y \\ax^2+2hxy+by^2=1\] We need to solve those 3 equations...

OpenStudy (freckles):

solve this system of 3 equations*